# SongBloom: Coherent Song Generation via Interleaved Autoregressive Sketching and Diffusion Refinement

### v2

@@ 1 Introduction
[…]
@@ 4.1 Task Formulation
MusicLM [23] proposes a two-stage framework that leverages semantic tokens, referred to as sketch, extracted from self-supervised learning (SSL) models such as MERT [24] as intermediate representations. In the first stage, a LM is employed to generate the sketch sequence from scratch, as its variable length and discrete nature make it well-suited for language modeling approaches, which have demonstrated strong scalability. In the second stage, either diffusion or LM subsequently generates acoustic tokens in a coarse-to-fine manner. The overall generation process can be formulated as:
@@ 4.3.3 Non-Autoregressive Latent Diffusion
Here, $t$ denotes the diffusion time step. $\textbf{z}^{i}_{t}$ represents the latent variables at time $t$ for the $i$-th patch, where $\textbf{z}^{i}_{0}=a_{(iP:(i+1)P]}$ and $\textbf{z}^{i}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ The variable $h_{i}$ is the hidden vector inherited from the first stage. $s_{(iP:(i+1)P]}$ refers to the sketch tokens for the current patch, and $\textbf{z}_{0}^{i-1}$ refers to the acoustic latents from the preceding patch.
@@
$\mathcal{L}_{\text{flow}}=\mathbb{E}_{t\sim\pi_{t},\textbf{z}^{i}\sim p(a)}\left[||v_{\phi}(t,\textbf{z}^{i}_{t}\mid\cdot)-(\textbf{z}^{i}_{1}-\textbf{z}^{i}_{0})||^{2}\right].$ (9)
The total training objective combines the sketch generation loss and the diffusion loss:
@@ 5.1 Dataset
We use a large-scale song dataset totaling 100K hours, including both Chinese and English songs, to train our model. The original lyrics are of relatively low quality, prompting the use of a dedicated data cleaning pipeline. First, we separate the vocal and instrumental tracks using the Demucs model [32]. Then we use the WhisperX toolkit [33] to filter out mismatched lyric sentences and recover missing words. The time boundaries of lyrics are also refined based on the results of automatic speech recognition (ASR). Additionally, structural information is extracted from the original tracks using the structure detector proposed by Kim and Nam [34].
@@ 5.2 Configuration & Training Setup
In our experiments, the codebook size for semantic tokens is set to 16,384, at a frame rate of 25. The stable-audio-vae [35] is adopted as the implementation of our waveform autoencoder. Minor modifications are made to its hyperparameters to ensure they have the same frame rate, thereby facilitating synchronized sketches and latents.
@@
All models are trained using the AdamW optimizer [38] with a learning rate of 1e-4. A cosine learning rate scheduler [39] with 2000 warm-up steps is employed to stabilize early training. Each model is trained for approximately 150K steps with a batch size of 128. The DeepSpeed strategy [40] is adopted to support efficient large-scale training. For inference, both stages share a classifier-free guidance coefficient of 1.5. Next-token prediction is performed using top-k sampling with $k=200$ and a temperature of 0.9. The diffusion process employs the standard Euler ODE solver with 36 diffusion steps.
@@ 5.3 Evaluation Metrics
We evaluate the proposed models using the following objective metrics: (1) Phoneme Error Rate (PER), computed based on the separated vocal tracks and the corresponding lyrics; (2) MuLan Cycle Consistency (MCC), which measures the cosine similarity of MuLan [41] embeddings between generated samples and reference audio or textual descriptions; (3) Fréchet Audio Distance (FAD)[42], which quantifies the distributional similarity between generated samples and real song clips; and (4) Structural Error Rate (SER), which measures the mismatch between detected structural patterns and the target lyric structure. The Dynamic Time Warping (DTW) algorithm [43] is first employed to obtain the optimal temporal alignment, and then we calculate the proportion of error duration. Additionally, we leverage the Audiobox-Aesthetic [44] to assess musical aesthetics, including content enjoyment (CE), content usefulness (CU), production complexity (PC), and production quality (PQ).
@@ 6 Results
Table 2: Objective evaluation results across all models. The upper section reports the performance of commercial platforms, while the lower section presents results for open-source baselines as well as our proposed models.
@@ 6.1 Full-Length Lyric-to-Song Generation
Based on the SongBloom-full model, we further fine-tuned it on synthesized data generated by Suno for an additional 1,000 steps, denoted as SongBloom-full-ft. This fine-tuning process enhances the model’s ability to emulate Suno-style compositions, leading to improved performance on lyrics with similar stylistic patterns. For the ACE-step baseline, the evaluation samples are generated via their HuggingFace demo page [^4] https://huggingface.co/spaces/ACE-Step/ACE-Step.
@@ 6.3 Effect of Hyper-Parameters
Figure 3 illustrates the impact of various patch sizes, where "+$\infty$+ ∞" denotes separating the generation of sketch and acoustic sequence. For speech-oriented autoregressive diffusion models [7, 19], a small patch size is essential, as larger patch sizes make the model difficult to converge. For SongBloom, since the sketch tokens have served as the CoT of hidden vectors and guide the diffusion directly, a small batch size is no longer the optimal choice. Although smaller patch sizes provide more acoustic information for sketch generation, thereby improving sketch token accuracy, they also hinder the fluency of acoustic latents during the diffusion stage, as the window for preceding contexts becomes smaller. Figure 3 illustrates the relationship between performance and diffusion steps during inference. The RTF increases proportionally with the number of diffusion steps, while near-optimal generation performance is achieved in as few as 10 steps, indicating that the inference process can be further accelerated.

### v4

@@ 1 Introduction
SongGen [9] LM 1.1B 30s 16
SongEditor [5] LM + Diff. 0.7B + 1B 120s 44.1
DiffRhythm-base [3] Diff. 1.1B 95s 44.1
DiffRhythm-full [3] Diff. 1.1B 285s 44.1
YuE [4] LM + LM 7B + 1B 300s 44.1
SongBloom-tiny LM & Diff. 1.3B 60s 48
SongBloom-full LM & Diff. 2B 150s 48
@@ 4.1 Task Formulation
MusicLM [23] proposes a two-stage framework that leverages semantic tokens, referred to as sketch, extracted from self-supervised learning (SSL) models such as MERT [24] as intermediate representations. In the first stage, an LM is employed to generate the sketch sequence from scratch, as its variable length and discrete nature make it well-suited for language modeling approaches, which have demonstrated strong scalability. In the second stage, either diffusion or LM subsequently generates acoustic tokens in a coarse-to-fine manner. The overall generation process can be formulated as:
@@ 4.3.3 Non-Autoregressive Latent Diffusion
Here, $t$ denotes the diffusion time step. In our experiments, we sample $t$ from a logit-normal distribution $\pi_{t}$ following Esser et al. [32]. $\textbf{z}^{i}_{t}$ represents the latent variables at time $t$ for the $i$-th patch, where $\textbf{z}^{i}_{0}=a_{(iP:(i+1)P]}$ and $\textbf{z}^{i}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ . The variable $h_{i}$ is the hidden vector inherited from the first stage. $s_{(iP:(i+1)P]}$ refers to the sketch tokens for the current patch, and $\textbf{z}_{0}^{i-1}$ refers to the acoustic latents from the preceding patch.
@@
$\mathcal{L}_{\text{flow}}=\mathbb{E}_{t\sim\pi_{t},\textbf{z}^{i}}\left[||v_{\phi}(t,\textbf{z}^{i}_{t}\mid\cdot)-(\textbf{z}^{i}_{1}-\textbf{z}^{i}_{0})||^{2}\right],$ (9)
where $\pi_{t}$ denotes the distribution over diffusion steps. The total training objective combines the sketch generation loss and the diffusion loss:
@@ 5.1 Dataset
We use a large-scale song dataset totaling 100K hours, including both Chinese and English songs, to train our model. The original lyrics are of relatively low quality, prompting the use of a dedicated data cleaning pipeline. First, we separate the vocal and instrumental tracks using the Demucs model [33]. Then we use the WhisperX toolkit [34] to filter out mismatched lyric sentences and recover missing words. The time boundaries of lyrics are also refined based on the results of automatic speech recognition (ASR). Additionally, structural information is extracted from the original tracks using the structure analyzer proposed by Kim and Nam [35].
@@ 5.2 Configuration & Training Setup
In our experiments, the codebook size for semantic tokens is set to 16384, at a frame rate of 25. The stable-audio-vae [36] is adopted as the implementation of our waveform autoencoder. Minor modifications are made to its hyperparameters to ensure they have the same frame rate, thereby facilitating synchronized sketches and latents.
@@
All models are trained using the AdamW optimizer [39] with a learning rate of 1e-4. A cosine learning rate scheduler [40] with 2000 warm-up steps is employed to stabilize early training. Each model is trained for approximately 150K steps with a batch size of 128. The DeepSpeed strategy [41] is adopted to support efficient training. For inference, both stages share a classifier-free guidance coefficient of 1.5. Next-token prediction is performed using top-k sampling with $k=200$ and a temperature of 0.9. The diffusion process employs the Euler ODE solver with 36 diffusion steps.
@@ 5.3 Evaluation Metrics
We evaluate the proposed models using the following objective metrics: (1) Phoneme Error Rate (PER), computed based on the separated vocal tracks and the corresponding lyrics; (2) MuLan Cycle Consistency (MCC), which measures the cosine similarity of MuLan [42] embeddings between generated samples and reference audio or textual descriptions; (3) Fréchet Audio Distance (FAD)[43], which quantifies the distributional similarity between generated samples and real songs from which the clips are intercepted; and (4) Structural Error Rate (SER), which measures the mismatch between detected structural patterns and the target lyric structure. The Dynamic Time Warping (DTW) algorithm [44] is first employed to obtain the optimal temporal alignment, and then we calculate the proportion of error duration. Additionally, we leverage the Audiobox-Aesthetic [45] to assess musical aesthetics, including content enjoyment (CE), content usefulness (CU), production complexity (PC), and production quality (PQ).
@@ 6 Results
Table 2: Objective evaluation results across all models. The upper section reports the performance of commercial platforms (valid until June 23, 2025), while the lower section presents results for open-source baselines as well as our proposed models.
@@ 6.1 Full-Length Lyric-to-Song Generation
Based on the SongBloom-full model, we fine-tuned it for 1,000 additional steps on synthesized data with a clear alternating structure of verses and choruses, denoted as SongBloom-full-ft. This fine-tuning process enhances the model’s ability to emulate synthesized lyric compositions, leading to improved performance on lyrics with similar stylistic patterns.
@@ 6.3 Effect of Hyper-Parameters
Figure 3 illustrates the impact of various patch sizes, where $\infty"$ denotes separating the generation of sketch and acoustic sequence. For speech-oriented autoregressive diffusion models [7, 19], a small patch size is essential, as larger patch sizes make the model difficult to converge. For SongBloom, since the sketch tokens have served as the CoT of hidden vectors and guide the diffusion directly, a small patch size is no longer the optimal choice. Although smaller patch sizes provide more acoustic information for sketch generation, thereby improving sketch token accuracy, they also hinder the fluency of acoustic latents during the diffusion stage, as the window for preceding contexts becomes smaller. Figure 3 illustrates the relationship between performance and diffusion steps during inference. The RTF increases proportionally with the number of diffusion steps, while near-optimal generation performance is achieved in as few as 10 steps, indicating that the inference process can be further accelerated.
@@ 7 Discussions and Limitations
8 Acknowledgement
This work is supported by National Natural Science Foundation of China (Grant No. 62271432), Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No. ZDSYS20230626091302006), Program for Guangdong Introducing Innovative and Entrepreneurial Teams, Grant No. 2023ZT10X044.